OpenStudy (anonymous):

Factor completely & show your work: 49m(squared) - 14m + 1

OpenStudy (lgbasallote):

\[49m^2 - 14m +1\] i can break down the middle term into \[49m^2 - 7m -7m + 1\] does that give you any ideas??

OpenStudy (anonymous):

ok so lets see if i can figure this out

OpenStudy (lgbasallote):

you can do it!!

OpenStudy (anonymous):

we can find its roots fro here

OpenStudy (anonymous):

7(7m*7m)-7(2m)+1

OpenStudy (lgbasallote):

what did you do?

OpenStudy (anonymous):

im not sure.... trying to read what the teacher did in school

OpenStudy (lgbasallote):

hmmm if i make it \[ (49m^2 - 7m) + (-7m + 1)\] does that help then?

OpenStudy (anonymous):

Have you heard of product, formula method of factoring?

OpenStudy (anonymous):

This is the example my teacher gave. Factor completely. .... (x+3y)(x-3y) Hint: same #'s different signs)..... x(x-3y)3y(x-3y) then x(squared)-3xy+3xy-9y. combine like terms so the 3xy cancel out. x(squared)-9y(squared) is the final answer. I'm not sure how to apply this to the problem

OpenStudy (anonymous):

(49m2−7m)+(−7m+1) This seems quite similar to x(x-3y)3y(x-3y) doesn't it? Only in your teacher's example, x and +3 have been factored out. Perhaps you can try and do that to this (49m2−7m)+(−7m+1)

OpenStudy (anonymous):

ok let me try

OpenStudy (anonymous):

I don't get it. I'm confused

OpenStudy (anonymous):

what do \[49m^2 \] and -7m have in common? 7m, so factor out 7m on the first one. to get 7m(7m-1) + (-7m+1) can you go on from there?

OpenStudy (lgbasallote):

\[(49m^2 - 7m) + (-7m+1)\] 49m^2 and -7m both have common factors of 7m so i can take it out \[7m(7m - 1) + (-7m +1)\] find a term that you can factor out of -7m + 1 to make it look like 7m - 1

OpenStudy (anonymous):

hold on please

OpenStudy (anonymous):

something is confusing me and im trying to figure out

OpenStudy (anonymous):

Hint: -7m+1 almost looks like 7m-1 apart from the difference in signs. So what would you factor out of -7m+1 to make it similar to 7m-1?

OpenStudy (lgbasallote):

i'll let you in a secret formua you WILL SURELY use a lot of times ;D \[(a-b) = -(-a + b) = -(b - a)\]

OpenStudy (anonymous):

which looks complicated. I think I am just not going to do this homework and let my teacher now that I need her to explain to me step by step. I appreciate all the help. It's just that I am extremely confused

OpenStudy (anonymous):

I'll explain it quickly using a different example. Take \[x ^{2}+2x-3\] If we want to factor this out we look at the coefficient of \[x ^{2}\], this is the number before \[x ^{2}\] , in this case it is one. If we had \[2x ^{2}\] the coefficient would be 2. So anyway, we multiply the coefficient of \[x^2\] which is one to -3, and then we say the prod=-3, then we look at 2x, the coefficient is 2x. So we say sum=+2 Now what sum and product(prod) mean here, is we need to find two numbers such that their product is -3 and their sum is +2. In the equation I gave, two numbers that would retriceup to +2 and have a product of -3 would be +3 and -1 because 3-1=2 and -1 *3=-3 So then we write the equation as \[x^2 -x + 3x -3\] notice that I put like terms together, x and x^2 and 3x and -3. So from here we factor, since \[x(x-1)+3(x-1)\] so we factor out the common terms and make it look like this. You would want to factor out the first terms to look like the second one so that you have for example x-1 twice in my example. Notice that x-1 repeats itself? So we write this equation as \[(x+3)(x-1)\] does this help?

OpenStudy (anonymous):

meant add up to* not retriceup